-(b-1)+12=-26 how do i determine what b is
Found 2 solutions by tutorcecilia, truman:
Answer by tutorcecilia(2152)

-(b-1)+12=-26 [Solve for b]
-b+1+12=-26 [Simplify]
-b+13=-26
-b+13-13=-26-13
-b=-39
-b/-1=-39/-1
b=39
.
Check by plugging (b=39) back into the original equation.

Answer by truman(1)

-b+1+12=-26
-b+13=-26
-b+13-13=-26-13
-b=-39
b=39